# A Sensitivity Analysis Framework for Quantifying Confidence in Decisions in the Presence of Data Uncertainty

### v1

@@ 1 Introduction
The CID analysis framework has features in common with the types of sensitivity analyses often done for causal inference (Fogarty, 2023) and for imputation of missing data. In causal inference, the goal of a sensitivity analysis typically is to determine how strong the effect of unmeasured confounding must be to nullify an apparent causal relationship (Rosenbaum, 2020; VanderWeele & Ding, 2017; Ding & VanderWeele, 2016; Schwartz et al., 2012; Mathur, 2023). With missing data, the goal of a sensitivity analysis typically is to assess how much estimates change under different specifications of missing not at random mechanisms (Little & Rubin, 2019; Linero & Daniels, 2018; Schifeling et al., 2015). Liublinska & Rubin (2014) propose tipping point analysis, which offers a visualization of how the p-value for test of no treatment effect changes under increasingly serious departures from missing at random assumptions. Our work contributes to these approaches by turning the focus explicitly on the impacts of assumptions on ultimate decisions as opposed to statistical estimates alone.
@@
The remainder of this article is organized as follows. In Section 2, we outline the CID framework. We do in the context of potential measurement error in the input to a predictive model for presidential election outcomes. In Section 3, we illustrate a a CID analysis with missing data. Here, the analyst considers how strong a missing not at random (MNAR) mechanism must be to change their decision compared to if the data were missing at random (MAR) and handled as such. The context is the percentage of children in North Carolina exposed to high levels of lead. We conclude in Section 4 with a discussion and avenues for future directions.
@@ 2 The CID Analysis Framework Illustrated Via Measurement Error in Model Inputs
We estimate the Bread and Peace model using the data provided in Gelman et al. (2021), displayed in Figure 1. The data cover every presidential election from 1952 to 2012. The data include the weighted average growth rate in real income over the previous presidential term used in Hibbs (2000) along with the vote share won by the incumbent’s party. The estimated model is
@@
Using (1), suppose that we seek to predict vote share in the 2024 presidential election from weighted average growth in real personal income throughout the previous four years, which we estimate with data from the Saint Louis Federal Reserve. In the 2024 election cycle, this input corresponds to $x=-0.728\%.$ Thus, the predicted vote share for the incumbent’s party (Democrat) in 2024 is 44.0% with a 95% prediction interval of (36.6%, 48.4%). The prediction interval clearly suggests that the challenger is favored to win (which indeed occurred). This is the decision reached from the analysis.
@@
When predicting elections, analysts can make one of three types of decisions: candidate A will win, candidate B will win, or it is not possible to predict a clear winner. We construct a decision rule as a map $\varphi(\ell,u)$ that takes the prediction interval $(\ell,u)$ for the vote share to a decision as follows,
@@
The rule in (2) decides the incumbent wins the election when the lower and upper bounds of the prediction interval are both above 50%, decides it is unclear who wins when the prediction interval overlaps with 50%, and decides the challenger wins when the bounds are both below 50%.
We are interested in the impact of measurement error in the input to (1) on our confidence in the decision in (2). To assess this impact, we perturb $x$ by a parameter $t$ corresponding to additive measurement error. For example, a value of $t=0.5$ corresponds to an analyst suspecting that the true value of the growth in personal income is 0.5 percentage points higher than the value $x$ used to make the prediction. The value $t=0$, which for convenience we label as $t_{0}$, corresponds to no measurement error. Absent specific information about the nature of the measurement error, we use $t_{0}$ for the baseline prediction from the Bread and Peace model. We then compute predicted vote share and associated prediction intervals for several plausible values of the true growth, $x+t$. We refer to this perturbation and computation step as “turning the knob” on the amount of measurement error. More extreme values of $t$ correspond to more extreme measurement error. For each of these perturbed inputs, we compute the CID metric.
@@
One can use $D_{t}$ as a CID metric. In some cases, however, analysts may want a more nuanced metric that reflects changes in the prediction intervals as one turns the knob for $t$. With this in mind, we illustrate a CID measure that assigns greater confidence in the decision when the prediction interval for some $t\neq t_{0}$ overlaps substantially with the prediction interval for $t_{0}$, and it assigns lesser confidence in the decision when the intervals based on $t$ and $t_{0}$ have little or no overlap. To encode this, we incorporate the confidence interval overlap metric of Karr et al. (2006). For any $t\neq t_{0}$, let $(L_{t},U_{t})$ and $(L_{t_{0}},U_{t_{0}})$ be the prediction intervals using $t$ and $t_{0}$, respectively, in (1). Let $(L_{\mathrm{over}},U_{\mathrm{over}})$ represent their intersection, defined respectively as the minimum and maximum values in the set, $\{b:b\geq L_{t_{0}},b\geq L_{t},b\leq U_{t_{0}},b\leq U_{t}\}$. When the set is empty, we let $L_{\mathrm{over}}=U_{\mathrm{over}}$. The overlap in the prediction intervals derived from the analysis with $t_{0}$ and the analysis with $t$ is
@@
We now use $\mathrm{CID}(t)$ to assess our confidence in the 2024 presidential election prediction from the Bread and Peace model. Figure 2 displays the values of the metric for $-4\leq t\leq 4$, assuming no measurement error as the reference analysis. This corresponds to a possible overestimate or underestimate in the real weighted average income growth of up to 4 percentage points. As evident in the top panel, once the measurement error $t>0.88$, so that the wage growth is truly at least 0.172, the decision changes, first to unclear winner at $t=0.88$ and then to the incumbent at $t=2.62$. For $-4\leq t\leq 0.88$, the ultimate decision to predict the challenger as the winner does not change. That is, the CID metric does not drop to zero over this range of $t$. The prediction interval based on $t_{0}$, i.e., no measurement error, overlaps substantially with the prediction intervals for many values of $t$; for example, $J_{t}\geq 0.5$ for $-2\leq t\leq 1.2$. Note that when $0.88\leq t\leq 1.2$, we have $J_{t}>0.5$ but the decision itself changes to an unclear winner, making $D_{t}=\mathrm{CID}(t)=0$ for $t$ in this range.
Analysts can use Figure 2 to visualize how quickly the confidence in decision changes with increasing measurement error. To utilize this figure most effectively, analysts need to determine what range of measurement errors is plausible. In other words, how far is too far to turn the knob for $t$? Ultimately, specifying plausible ranges of $t$ for a specific CID analysis should be based on domain expertise. For this problem, we define a plausible range of error by considering the widest range of growth or loss in the four year window for which the weighted average was computed. In the case of the 2024 election, using election day in 2019 as a starting point, the largest loss was 1.363% and the gain never crosses 0%. This corresponds to a plausible region for $t$ of $(-0.635,0.728)$. As evident in Figure 2, the CID metric does not change substantially within this plausible range of measurement error. As a result, the analyst can conclude with reasonable confidence that their decision with the Bread and Peace model is insensitive to plausible measurement error in $x$.
@@ 3 Example of CID Analysis with Missing Data
We now illustrate a CID analysis where the data imperfection is missing data. Typically, analysts presume that values are MAR. This may be partly because standard statistical methods and software are designed for missingness at random. Regardless, the analyst may be concerned that a decision reached under MAR assumptions could change if the data are MNAR, and hence seek to assess their confidence in the decision under departures from a default MAR modeling approach. We illustrate such analyses using data on lead exposure among children in North Carolina, as we now describe.
Exposure to high levels of lead can have adverse effects on children’s cognitive development (Bellinger et al., 1992; Miranda et al., 2009; Bravo et al., 2022). Thus, public health decision-makers may choose to intervene when the percentage of children exposed to high levels of lead is deemed unacceptable. To determine these percentages, one approach is to utilize large administrative health databases, which may contain tens of thousands of tested children. However, there can be selection mechanisms into these databases; for example, children who are suspected to have high levels of lead exposure may be more likely to be tested for that exposure. Thus, an analyst computing lead exposure rates from administrative data should evaluate their confidence in their ultimate decision about the percentage of children at unacceptably high levels and any resulting policy intervention.
@@
Table 1: Empirical distribution of lead exposures among the children in the North Carolina lead data. Smaller values indicate lesser amounts of lead exposure. Estimates taken from Feldman et al. (2024).
For our analysis, we use administrative data from the state of North Carolina for children born between 2003 and 2005. These data were used by Feldman et al. (2024), among others, to assess correlates of lead exposure and educational outcomes. The data comprise approximately $n=110,000$ children in the desired age range who have recorded, integer-valued blood lead levels. Table 1 displays the empirical percentages. According to the Kids Count Data Center (2020), approximately $N=400,000$ children were born in North Carolina between 2003 and 2005. Hence, considering the $n$ children as part of the population, we are missing lead measurements for approximately $290,000$ children. Here, we presume no other information is available on these children. This is actually the case for us as we do not have access to the record-level data, which are confidential. We discuss extensions to CID analysis with multivariate data in Section 4.
@@
As the strategy for handling the missing data, we presume the analyst follows typical practice and applies a missing (completely) at random model for multiple imputation. We assume that lead levels follow a multinomial distribution with support on the integers from 1 to 10. We use uniform prior distributions on the multinomial probabilities. Thus, to make any single completed dataset comprising the $n$ observed and $N-n$ imputed lead values, under MAR we first draw a value of the $10$-dimensional multinomial probability vector $\mathbf{p}\sim\mathrm{Dirichlet}(1+n_{1},\dots,1+n_{10})$. Given the sampled value of $\mathbf{p}$, we draw a value $y_{i}\sim\mathrm{Multinomial}(\mathbf{p})$ for all units with $r_{i}=1.$ We repeat this procedure independently $M=5$ times to create $M$ completed populations, $D^{(1)},\dots,D^{(M)}$. The analyst can use the combining rules from Rubin (1987) to make inferences about any estimands, including ones intended to inform a policy-maker’s decision to intervene or not.
@@ 3.1 Step 1: Specify decision rule
To construct an illustrative decision rule, we presume that a governmental agency provides intervention services when 20% or more of the children in the state have lead exposure levels beyond 3. This decision rule is motivated by targets for lead exposure rates in the Healthy People 2030 guidelines (Office of Disease Prevention and Health Promotion, 2025). In practice, of course, public policy-makers could select the level or percentage that they deem most suitable to trigger interventions for the specific matter at hand. Our decision rule is a map $\varphi(\hat{\theta})$ that takes the estimated percentage $\hat{\theta}$ of children with lead level above 3 to a decision as follows.
@@ 3.2 Step 2: Estimate under departures from missing at random
We represent a MNAR mechanism by adding the vector $\boldsymbol{\alpha}_{t}=(\alpha_{1},\dots,\alpha_{10})\neq\bf{0}$ to $\boldsymbol{\beta}_{0}$. Let $\boldsymbol{\beta}_{t}=\boldsymbol{\beta}_{0}+\boldsymbol{\alpha}_{t}$. We use $\boldsymbol{\beta}_{t}$ to define the probabilities for the missing data imputation. That is, we impute each $y_{i}$ with $r_{i}=1$ using the multinomial distribution with probabilities
@@ 3.3 Step 3: Construct CID metric
We next construct the CID metric. We do so in four parts for this analysis. Let $D_{t}=1$ when the decision to intervene or not as determined by (5) is the same under the MAR mechanism ($t=t_{0}=0$) and under the MNAR mechanism encoded using $t\neq 0$. When $D_{t}=1$, we let $\mathrm{CID}(t)=1$.
When $D_{t}=0,$ the value of $\mathrm{CID}(t)$ depends on the reference decision. If $\hat{\theta}\leq 0.20$ and $\hat{\theta}_{t}>0.20$, the analyst who assumes lead values are MAR underestimates the proportion of children with concerningly high lead levels. In this case, we set
$\mathrm{CID}(t)=1-\frac{\hat{\theta}_{t}-0.20}{\theta_{\max}-0.20}.$ (10)
Here, $\theta_{\max}$ is the percentage of children whose levels exceed 3 if every single child with $r_{i}=1$ has a lead level above 3. Essentially, (10) encodes a varying cost of making a wrong decision. The maximum cost equals one, which occurs when the true rate $\hat{\theta}_{t}$ is the worst it possibly can be (i.e., everyone missing a lead measurement exceeds the threshold of 3). The minimum cost equals zero, which occurs when $\hat{\theta}_{t}$ is right at the threshold.
Likewise, when $\hat{\theta}>0.20$ and $\hat{\theta}_{t}\leq 0.20$, the analyst using a MAR assumption overestimates the percentage of children whose lead levels exceed 3. In this case, we encode a cost to making the wrong decision, as the policy-maker intervenes when not necessary. We set
$\mathrm{CID}(t)=1-\frac{0.20-\hat{\theta}_{t}}{0.20-\theta_{\min}}.$ (11)
We determine $\theta_{\min}$ by imputing all $y_{i}\leq 3$ for children with $r_{i}=1$. We note that analysts could choose other cost functions, for example, to encode larger costs for underestimating high lead exposure rates than overestimating them. These are specific to the context at hand; for example, they could be based on valuations of the trade off in risks to child health and financial expenditures.

### v2

@@ 1 Introduction
The CID analysis framework has features in common with the types of sensitivity analyses often done for causal inference (Fogarty, 2023) and for imputation of missing data. In causal inference, the goal of a sensitivity analysis typically is to determine how strong the effect of unmeasured confounding must be to nullify an apparent causal relationship (Rosenbaum, 2020; VanderWeele & Ding, 2017; Ding & VanderWeele, 2016; Schwartz et al., 2012; Mathur, 2023). With missing data, the goal of a sensitivity analysis typically is to assess how much estimates change under different specifications of missing not at random mechanisms (Little & Rubin, 2019; Linero & Daniels, 2018; Schifeling et al., 2015). Liublinska & Rubin (2014) propose tipping point analysis, which offers a visualization of how the $p$-value for a test of no treatment effect changes under increasingly serious departures from missing at random assumptions. Our work contributes to these approaches by turning the focus explicitly on the impacts of assumptions on ultimate decisions as opposed to statistical estimates alone.
@@
The remainder of this article is organized as follows. In Section 2, we outline the CID framework. We do in the context of potential measurement error in the input to a predictive model for presidential election outcomes. In Section 3, we illustrate a a CID analysis with missing data. Here, the analyst considers costs of using the estimate from a missing at random (MAR) analysis when in fact the observed data are missing not at random (MNAR) to varying degrees. The context is the percentage of children in North Carolina exposed to high levels of lead. We conclude in Section 5 with a discussion and avenues for future directions. R code to reproduce the results is available at: https://github.com/adway/CID-analysis.
@@ 2 The CID Analysis Framework Illustrated Via Measurement Error in Model Inputs
We estimate the Bread and Peace model using the data provided in Gelman et al. (2021), displayed in Figure 1. The data cover every presidential election from 1952 to 2012. The data include the weighted average growth rate in real income over the previous presidential term along with the vote share won by the incumbent’s party. The estimated model is
@@
Using (1), suppose that we seek to predict vote share in the 2024 presidential election from weighted average growth in real personal income throughout the previous four years, which we estimate with data from the Saint Louis Federal Reserve. In the 2024 election cycle, this input corresponds to $x=-0.728\%.$ Thus, the predicted vote share for the incumbent’s party (Democrat) in 2024 is 44.0% with a 95% prediction interval of (39.6%, 48.4%). The prediction interval clearly suggests that the challenger is favored to win (which indeed occurred). This is the decision reached from the analysis.
@@
When predicting elections, analysts can make one of three types of decisions: candidate A will win, candidate B will win, or it is not possible to predict a clear winner. We construct a decision rule as a map $\varphi(\ell,u)$ that takes the prediction interval $(\ell,u)$ for the vote share ($\times 100$), where $\ell$ and $u$ are the lower and upper bounds of the interval, respectively, to a decision as follows,
@@
The rule in (2) decides the incumbent wins the election when the lower and upper bounds of the prediction interval are both above 50%, decides it is unclear who wins when the prediction interval includes 50%, and decides the challenger wins when the bounds are both below 50%.
We are interested in the impact of measurement error in the input to (1) on our confidence in the decision in (2). To assess this impact, we perturb $x$ by a parameter $t$ corresponding to additive measurement error. For example, a value of $t=0.5$ corresponds to an analyst suspecting that the true value of the growth in personal income is 0.5 percentage points higher than the value $x$ used to make the prediction. The value $t=0$, which for convenience we label as $t_{0}$, corresponds to no measurement error. Absent specific information about the nature of the measurement error, we use $t_{0}$ for the baseline prediction from the Bread and Peace model. We then compute predicted vote share and associated prediction intervals for several plausible values of the true growth, $x+t$. For example, with $t=0.5$, the predicted vote share is 46.248+3.061(0.728+0.5)46.248+3.061(-0.728+0.5). We refer to this perturbation and computation step as “turning the knob” on the amount of measurement error. More extreme values of $t$ correspond to more extreme measurement error. For each of these perturbed inputs, we compute the CID metric.
@@
One can use $D_{t}$ as a CID metric. In some cases, however, analysts may want a more nuanced metric that reflects changes in the prediction intervals as one turns the knob for $t$. With this in mind, we construct a CID measure that assigns greater confidence in the decision when the prediction interval for some $t\neq t_{0}$ overlaps substantially with the prediction interval for $t_{0}$, and it assigns lesser confidence in the decision when the intervals based on $t$ and $t_{0}$ have little or no overlap. To encode this, we incorporate the confidence interval overlap metric of Karr et al. (2006). For any $t\neq t_{0}$, let $(L_{t},U_{t})$ and $(L_{t_{0}},U_{t_{0}})$ be the prediction intervals using $t$ and $t_{0}$, respectively, in (1). Let $(L_{\mathrm{over}},U_{\mathrm{over}})$ represent their intersection, defined respectively as the minimum and maximum values in the set, $\{b:b\geq L_{t_{0}},b\geq L_{t},b\leq U_{t_{0}},b\leq U_{t}\}$. When the set is empty, we let $L_{\mathrm{over}}=U_{\mathrm{over}}$. The overlap in the prediction intervals derived from the analysis with $t_{0}$ and the analysis with $t$ is
@@
We now use $\mathrm{CID}(t)$ to assess our confidence in the 2024 presidential election prediction from the Bread and Peace model. Figure 2 displays the values of the metric for $4\leq t\leq 4$, assuming no measurement error as the reference analysis. This corresponds to a possible overestimate or underestimate in the real weighted average income growth of up to 4 percentage points. As evident in the top panel, once the measurement error $t>0.88$, so that the wage growth is truly at least 0.172, the decision changes, first to unclear winner at $t=0.88$ and then to the incumbent at $t=2.62$. For $4\leq t\leq 0.88$, the ultimate decision to predict the challenger as the winner does not change. That is, the CID metric does not drop to zero over this range of $t$. The prediction interval based on $t_{0}$, i.e., no measurement error, overlaps substantially with the prediction intervals for many values of $t$; for example, $J_{t}\geq 0.5$ for $2\leq t\leq 1.2$. Note that when $0.88\leq t\leq 1.2$, we have $J_{t}>0.5$ but the decision itself changes to an unclear winner, making $D_{t}=\mathrm{CID}(t)=0$ for $t$ in this range.
Analysts can use Figure 2 to visualize how quickly the confidence in decision changes with increasing measurement error. To utilize this figure most effectively, analysts need to determine what range of measurement errors is plausible. In other words, how far is too far to turn the knob for $t$? Ultimately, specifying plausible ranges of $t$ for a specific CID analysis should be based on domain expertise. For this problem, we define a plausible range of error by considering the widest range of growth or loss in the four year window for
@@
which the weighted average was computed. In the case of the 2024 election, using election day in 2019 as a starting point, the largest loss was 1.363% and the gain never crosses 0%. This corresponds to a plausible region for $t$ of (0.635,0.728)(-0.635,0.728). As evident in Figure 2, the CID metric does not change substantially within this plausible range of measurement error. As a result, the analyst can conclude with reasonable confidence that their decision with the Bread and Peace model is insensitive to plausible measurement error in $x$.
Analysts can use CID analysis with other ways of accounting for potential measurement error. As one example, to make a prediction interval for the 2024 vote share, analysts can specify and average over a distribution of the measurement error for $x$, such as an additive measurement error model where $t\sim N(0,\tau^{2})$ for some selected variance $\tau^{2}.$ By averaging over this distribution, the analyst can obtain a single prediction interval with upper and lower limits that can be inputs to the CID metric. Analysts could repeat the analysis using different measurement error distributions, e.g., via different values of $\tau^{2}$, in which case $\tau^{2}$ becomes the knob to enable a CID analysis.
@@ 3 Example of CID Analysis with Missing Data
We now illustrate a CID analysis where the data imperfection is missing data. Typically, analysts presume that values are MAR. This may be partly because standard statistical methods and software are designed for missingness at random. Regardless, the analyst may be concerned that a decision reached under MAR assumptions could be incorrect if the data are MNAR and hence seek to assess their confidence in the decision under departures from a default MAR modeling approach. We illustrate such analyses using data on lead exposure among children in North Carolina, as we now describe.
Exposure to high levels of lead can have adverse effects on children’s cognitive development (Bellinger et al., 1992; Miranda et al., 2009; Bravo et al., 2022). Thus, public health decision-makers may choose to intervene when the percentage of children exposed to high levels of lead is deemed unacceptable. To determine these percentages, one approach is to utilize large administrative health databases, which may contain tens of thousands of tested children. However, there can be selection mechanisms into these databases; for example, children who are suspected to have high levels of lead exposure may be more likely to be tested for that exposure. Thus, an analyst computing lead exposure rates from administrative data should evaluate their confidence in their ultimate decision about the percentage of children at unacceptably high levels, as well as any resulting policy intervention.
@@
Table 1: Empirical distribution of lead exposures (micrograms per deciliter) among the children in the North Carolina lead data. Smaller values indicate lesser amounts of lead exposure. Estimates taken from Feldman et al. (2024). The blood lead measurements in the data are available only as integers; we do not have access to values with greater precision.
For our analysis, we use administrative data from the state of North Carolina for children born between 2003 and 2005. These data were used by Feldman et al. (2025+), among others, to assess correlates of lead exposure and educational outcomes. The data comprise approximately $n=110,000$ children in the desired age range who have recorded, integer-valued blood lead levels. Table 1 displays the empirical percentages. According to the Kids Count Data Center (2020), approximately $N=400,000$ children were born in North Carolina between 2003 and 2005. Hence, considering the $n$ children as part of the population, we are missing lead measurements for approximately $290,000$ children. Here, we presume no other information is available on these children. This is actually the case for us as we do not have access to the record-level data, which are confidential.
@@
As the strategy for handling the missing data, we presume the analyst follows typical practice and applies a missing (completely) at random model for multiple imputation. We assume that lead levels follow a multinomial distribution with support on the integers from 1 to 10. We use uniform prior distributions on the multinomial probabilities. Thus, to make any single completed dataset comprising the $n$ observed and NnN-n imputed lead values, under MAR we first draw a value of the $10$-dimensional multinomial probability vector $\mathbf{p}\sim\mathrm{Dirichlet}(1+n_{1},\dots,1+n_{10})$. Given the sampled value of $\mathbf{p}$, we draw a value $y_{i}\sim\mathrm{Multinomial}(\mathbf{p})$ for all units with $r_{i}=1.$ We repeat this procedure independently $M=5$ times to create $M$ completed populations, $D^{(1)},\dots,D^{(M)}$. The analyst can use the combining rules from Rubin (1987) to make inferences about any estimands, including ones intended to inform a policy-maker’s decision to intervene or not.
@@ 3.1 Step 1: Specify decision rule
[…]
@@ 3.2 Step 2: Estimate under departures from missing at random
We represent a MNAR mechanism by adding some vector $\boldsymbol{\alpha}_{t}=(\alpha_{1},\dots,\alpha_{10})\neq\bf{0}$ to $\boldsymbol{\beta}_{0}$; we examine two choices for $\boldsymbol{\alpha}_{t}$ below. Let $\boldsymbol{\beta}_{t}=\boldsymbol{\beta}_{0}+\boldsymbol{\alpha}_{t}$. We use $\boldsymbol{\beta}_{t}$ to define the probabilities for the missing data imputation. That is, we impute each $y_{i}$ with $r_{i}=1$ using the multinomial distribution with probabilities
@@ 3.3 Step 3: Construct CID metric
We next construct the CID metric. Let $D_{t}=1$ when the decision to intervene or not as determined by (5) is the same under the MAR mechanism ($t=t_{0}=0$) and under the MNAR mechanism encoded using $t\neq 0$. We use a $\mathrm{CID}(t)$ that depends on the value of $\hat{\theta}$ in the reference analysis—in our case, the estimate assuming MAR—as we now describe.
First consider when $\hat{\theta}\geq 0.20.$ In this case, the decision is to intervene. We presume the decision-makers intend to dedicate sufficient resources to the intervention to reduce the percentage of children with high lead exposure $\theta$ to the target of 20%. If the true $\theta$ happens to be less than $\hat{\theta}$, i.e., $\hat{\theta}$ is an overestimate, the decision-makers would spend resources on the intervention unnecessarily; this should be considered a negative in the CID metric. On the other hand, if $\theta\geq\hat{\theta}$, the decision makers would not meet the target; this too should be considered a negative in the CID metric. We encode these considerations using
$\mathrm{CID}(t)=I(\hat{\theta}>.20)\left(1-\frac{I(\hat{\theta}\geq\hat{\theta}_{t})\min(\hat{\theta}-0.20,\hat{\theta}-\hat{\theta}_{t})a+I(\hat{\theta}<\hat{\theta}_{t})(\hat{\theta}_{t}-\hat{\theta})b}{C}\right),$ (10)
where $I(q)=1$ when the condition $q$ is true and $I(q)=0$ otherwise. Here, $a$ and $b$ are cost parameters set by the analyst and decision-makers. Specifically, $a$ represents the cost of spending resources unnecessarily on an intervention, which is triggered by the overestimation of $\theta$. And, $b$ represents the cost of not intervening to reduce the fraction of children with high lead exposure to the target level of 20%. In our analysis, we interpret $a$ and $b$ as costs per 1% of children estimated to have high lead exposure. Additionally, $C$ is the maximum cost of inaccurate estimation. In (10), we set $C=\max((\hat{\theta}-.20)a,(\theta_{\mathrm{WC}}–\max(\hat{\theta},0.20))b)$, where $\theta_{\mathrm{WC}}$ is the worst case percentage of children whose levels exceed 3 if every single child with $r_{i}=1$ has a lead level above 3. Thus, the $\mathrm{CID}(t)=1$ when $\theta=\hat{\theta}$ and reduces to zero for the $\theta$ with the largest cost possible.
To understand this CID metric further, consider the value of (10) when $\hat{\theta}=0.25$ and the true percentage of children with high lead exposure is one of $\hat{\theta}_{t}\in\{0.15,0.22,0.27$}. For $\hat{\theta}_{t}=0.15$, the decision should be not to intervene, and the value of (10) reflects the $5a$ cost of that unnecessary intervention. For $\hat{\theta}_{t}=0.22$, the decision to intervene is right, but an additional $3a$ in cost was incurred unnecessarily. For $\hat{\theta}_{t}=0.27$, $D_{t}=1$ but an additional 2% of children have high lead exposure than estimated, leading to $2b$ in costs. These additional costs would be scaled by $C$ to facilitate interpretations and visualizations.
We also need to specify the CID metric when $\hat{\theta}<0.20$. In this case, the decision under the reference MAR analysis is not to intervene. The decision-maker incurs no cost for any value of $\theta$ that does not change the decision, i.e., for which $D_{t}=1$. However, when $\theta>.20$, $D_{t}=0$ and the decision maker should have dedicated sufficient resources for intervention to meet the target. We encode this by setting
$\mathrm{CID}(t)=I(\hat{\theta}<0.2)\left(1-\frac{(1-D_{t})(\hat{\theta}_{t}-0.20)b}{C}\right).$ (11)
To illustrate, suppose that $\hat{\theta}=0.18$. For $\hat{\theta}_{t}=0.19$, the CID metric in (11) encodes no loss, since not intervening is still the right decision to stay under the target. For $\hat{\theta}_{t}=.27,$ the decision-maker should have intervened, and the cost of missing 7% of kids above the 20% target is $7b$. This $7b$ cost is scaled by the maximum possible cost of $C$.
In our analysis, we set $a=b$, equating the cost of failing to provide enough resources for 1% of children to not spending $a$ in resources. Since $\hat{\theta}=.25>.20$, we use (10) with $C=.54$ as ${\theta}_{WC}=0.79$. Of course, this value of ${\theta}_{WC}$ is completely unrealistic; however, $C$ solely serves as a scaling factor for purposes of visualization. Analysts could set $(a,b)$ otherwise based on policy evaluations (Gould, 2009). More generally, the selection of $(a,b)$ is specific to the context at hand; for example, they could be based on valuations of the trade off in risks to child health and financial expenditures.